OpenStudy (anonymous):

The demand for fax machines as part of an electronics business has decreased from 20% in 2005 to 10% in 2010 due to other Internet-based options for the transfer of documents. Assuming that this decrease follows a linear model, what is the model?

OpenStudy (anonymous):

A linear model follows the equation y=mx+b. Then, you need to find m and b. To find m, the slope, take the difference in % over the difference in years. So, (10%-20%)/(2010-2005)=-10%/5years = -2% per year. Now, find b, the y intercept, by plugging in one set of points. You can use the 2010 at 10%. So, 10%=(-2% per year)*(2010)+b. Now, solve for b. b=10+4020 = 4030. Now plug m and b into y=mx+b